OpenStudy (anonymous):

The sum of the perimeters of an equilateral triangle and a square is 13. Find the dimensions of the triangle and the square that produce a minimum total area.

OpenStudy (anonymous):

wow... ok

OpenStudy (anonymous):

so du u know formula for area of square and area of triangle @akt_tippit ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sweet, so somehow we gotta make this a simultaneous eqn or a differential eqn... so Area of square = s^2, yeah? and Area of Triangle = 1/2 * base * ht, yeah? and equilateral triangle soo... gimme a sec

OpenStudy (anonymous):

Area of equilateral Triangle = \[A = \frac {\sqrt 3} 4 \times base^2\]

OpenStudy (anonymous):

agreed? @akt_tippit ?

OpenStudy (anonymous):

agreed!

OpenStudy (anonymous):

sweet, so now we hafta put in the perimeter side of this into the combined area equation, so we're only dealin with one equation that way... sooooooo

OpenStudy (anonymous):

perimeter of an equilateral triangle = base + base + base = 3b perimeter of a square = side + side + side + side = 4s so if perimeter = 13 (total) 3b + 4s = 13 still with me man?

OpenStudy (anonymous):

and combined area eqn is A square + Area triangle Total Area = s^2 + sqrt 3/4 * b^2

OpenStudy (anonymous):

so i think that's our 2 equations rite there...

OpenStudy (anonymous):

so we solve for b and then put that equation into the total area equation

OpenStudy (anonymous):

??

OpenStudy (anonymous):

yep, sounds like a plan man, u go ahead and i'll watch an see how it goes?

OpenStudy (anonymous):

jus lemme know if u want help with anything, k?

OpenStudy (anonymous):

okay, so when you plug b in it becomes Square-root3/4(13-4s/3)^2+s^2 then square root 3/4(169/9 - 16s^2/9) + s^2? (I don't think I squared it right though) then 169square-root3/36 -16square-root 3 s^2/36 +s^2? I think I messed up somewhere

OpenStudy (anonymous):

k heres what i got so far:

OpenStudy (anonymous):

so \[3b + 4s = 13\] which means \[3b = 13 - 4s\] so \[\large b = \frac {13 - 4s}3\] \[\large Total~Area = s^2 + (\frac {\sqrt 3}4 \times b^2)\] so if \[\large b = \frac {13 - 4s}3\] then \[\large Total~Area = s^2 + (\frac {\sqrt 3}4 \times (\frac {13 - 4s}3)^2)\] so the hard part: expand out the b^2 \[(\frac {13 - 4s}3)^2\] \[(\frac {13 - 4s}3)(\frac {13 - 4s}3)\] so use FOIL and u should get \[\frac {16s^2 - 104s +169}9\] now put that back into the Total Area Eqn \[\large Total~Area = s^2 + [\frac {\sqrt 3}4 \times (\frac {13 - 4s}3)^2]\] \[\large Total~Area = s^2 + [\frac {\sqrt 3}4 \times (\frac {16s^2 - 104s +169}9)]\] \[\large Total~Area = s^2 + [\frac {\sqrt 3}4 \times (\frac {16s^2 - 104s +169}9)]\]

OpenStudy (anonymous):

\[\large Total~Area = s^2 + [\frac {\sqrt 3}4 \times (\frac {16s^2 - 104s +169}9)]\] \[\large Total~Area = s^2 + [ (\frac {\sqrt 3 \times (16s^2 - 104s +169)}{4 \times 9})]\] \[\large Total~Area = s^2 + [ (\frac {\sqrt 3 \times (16s^2 - 104s +169)}{4 \times 9})]\] \[\large Total~Area = s^2 + [ \frac { 16s^2 - 104s +169}{\sqrt 3 \times 12}]\] wow... maybe we should have solved for s in the first eqn ;D, anyways, thats the equation hay now iwe differentiate to find zero

OpenStudy (anonymous):

so expanded form to put into y = ax^2 + bx + c \[\large A = \frac {4s^2}{3 \sqrt3} + s^2- \frac{26 s}{3 \sqrt3} + \frac {169}{12 \sqrt3}\] \[\large A = \frac {4+ 3 \sqrt3 }{3 \sqrt3}s^2 - \frac{26 s}{3 \sqrt3} + \frac {169}{12 \sqrt3}\] so that can be solved by using quadratic equation: \[\large y = ax^2 + bx + c\] \[\large y = \frac {4+ 3 \sqrt3 }{3 \sqrt3}x^2 - \frac{26 x}{3 \sqrt3} + \frac {169}{12 \sqrt3}\] us quadratic equation: \[\huge x = \frac {-b± \sqrt{b^2-4 ac}}{2a}\] now plug in your values for a, b and c above and you'll get x = 1.4136 ish so x is s so a square with a side length of 1.4136

OpenStudy (anonymous):

so use that value of s to work out b, so if u go back to 3b + 4s = 13 u can solve for b from here those r ur dimensions for min area @akt_tippit

OpenStudy (wolf1728):

Wow Jack, that was a LOT of work

OpenStudy (anonymous):

aww thanks wolf! OP ran away... was kinda sad for a bit, love feelin appreciated so cheers man!

OpenStudy (wolf1728):

okay and cheers to you Jack gee at least the OP could have left you a medal.

OpenStudy (anonymous):

all good, my bad, maybe i should have rounded heaps earlier, thos sqrt signs are kinda scary ;D

OpenStudy (wolf1728):

LOL well see ya around here no doubt

OpenStudy (anonymous):

4 sure man, slaters!

OpenStudy (anonymous):

How many hours did you spend on this thread?

OpenStudy (anonymous):

Oh nevermind I see the timestamp. ~1 hour.

OpenStudy (anonymous):

Good LaTeX skills right there. lol

OpenStudy (anonymous):

cheers man!

OpenStudy (anonymous):

jus hope OP comes back so is not 1 hr wasted xD

OpenStudy (anonymous):

That was awesome! I was so frustrated with it and you helped out so much! thanks Jack!!!